# Understanding and Supporting Formal Email Exchange by Answering AI-Generated Questions

### v1

@@ 6.2.1. Perceived Quality of the Email by Evaluators (H2)
In Fig. 6, H2 shows the results regarding the quality of the emails. The Cronbach’s Alpha of three items measuring the perceived quality of the email is $0.846$. After checking the data normality assumption with the Shapiro-Wilk test, the result of one-way repeated measure ANOVA showed that there was a significant difference in the perceived quality of the email across three conditions $(F[2,22]=9.1,p=0.001{\color[rgb]{0,0,0},\eta_{p}^{2}=0.45})$. Post-hoc analysis with Holm correction revealed that the perceived quality of the emails participants’ wrote in the QA-based condition was significantly higher compared to the No-AI $(t(11),p=0.005{\color[rgb]{0,0,0},d=1.21})$ condition. Thus, H2 was partially supported. The QA-based approach improved the quality of the email responses compared to No-AI.
@@ 6.3.1. Perceived Impression of Participants by Evaluators (H3-a)
The results of the perceived impression of the participants rated by another group of evaluators are shown in Fig. 6 H3-a. The two items assessing participants’ impression as email senders showed high internal consistency, with a Cronbach’s Alpha of $0.946$. After checking the data normality assumption with the Shapiro-Wilk test, the result of one-way repeated measure ANOVA showed that there was a significant difference in impression of the participants as an email sender across three conditions $(F[2,22]=5.9,p=0.009{\color[rgb]{0,0,0},\eta_{p}^{2}=0.35})$. Post-hoc analysis with Holm correction revealed that participants’ impression in the QA-based condition was not significantly higher compared to both the No-AI $(t(11),p=0.058{\color[rgb]{0,0,0},d=0.79})$ and the Prompt-based $(t(11),p=0.939{\color[rgb]{0,0,0},d=0.02})$ conditions. Thus, H3-a was not supported. The QA-based approach didn’t improve the impression of participants as email senders.
@@ 6.4.1. Participants’ Email-Replying Process (RQ1)
  1. (1)
    “In the QA-based condition, AI summarized key points through questions and highlighted relevant sections of the email body, which facilitated my understanding of the email and reduced my overall burden” [P10].
  2. (2)
    “In the QA-based condition, I could easily obtain the desired output even without the technical skills to create prompts” [P6].
  3. (3)
    “By saving the time needed to read the counterpart’s text, the psychological barrier to starting the task was lowered” [P5].
@@
  1. (1)
    “Since the AI prompted me with questions at the beginning, the mental effort required to start thinking about the task was eliminated, reducing the stress associated with initiating the work” [P10].
  2. (2)
    “By saving the time needed to read the counterpart’s text, the psychological barrier to starting the task was lowered” [P5].
@@ 6.4.3. Relationship between Participants and Their Counterpart (RQ3)
Participants shared differing views on how AI’s involvement affected their psychological distance from their counterparts. P2, P9, and P11 reported that the psychological distance they felt from the other person was directly related to the amount of effort they put in. Furthermore, P6 noted that “especially under QA-based condition, I barely thought about the counterpart because I only selected options to create responses”. In contrast, P8 reported that “compared to composing replies myself, using AI allowed me to create messages that left a better impression on my counterpart, which made the relationship feel closer” These results suggested that, on the one hand, AI’s mediation can potentially increase the psychological distance between senders and receivers. On the other hand, it can also diminish the perceived distance from the sender due to effective impression management. Thus, we conducted a field study to further clarify the impact of AI on interpersonal relationships.
@@ 9.1.2. Potential Reduction in Sense of Agency and Control
While the QA-based approach enhanced users’ efficiency, our studies also revealed a potential trade-off in users’ sense of agency and control (Sec. 6.1.76.4.18.1.3). Some participants reported a decreased sense of authorship, feeling more like editors than creators of their emails. This reduction in agency may be due to the diminished amount of text input required from the user, as the AI takes a more active role in content generation. Moreover, we found that a sense of agency influenced users’ preferences for future usage. Among those participants who still maintained their sense of agency, we found that they tended to actively review and modify the AI-generated content to reflect their personal style and intentions. This suggests that even when AI intervention is substantial, users can maintain a sense of authorship by actively engaging with and refining the AI’s suggestions. To optimize users’ level of agency, adapting the degree of AI intervention in the email construction process can be helpful. For instance, by adjusting the number and type of AI-generated questions or varying the levels of AI-generated suggestions (Fu et al., 2023), ranging from word-level to message-level.
@@ 10. Conclusion
In formal email communication, users are often required to read detailed (lengthy or complex) emails. Crafting appropriate responses to such emails is time-consuming and may lead to overlooked sender requests or delayed responses, causing communication issues. Thus, we propose QA-based approach, which leverages LLM-based question generation to help users create efficient and high-quality replies by generating multiple question-answer pairs related to the received email content. To examine the comprehensive impact of the QA-based approach on both email senders and recipients, we conducted controlled and field experiments using our prototype system, ResQ. Our findings demonstrate that structuring email content into question-answer pairs improves efficiency, reduces cognitive load, and lowers the barriers to initiating responses. Additionally, this approach enhances email quality and may leave a better impression on recipients. However, our findings also revealed challenges, including a potential reduction in user agency and an increased psychological distance in communication. These trade-offs emphasize the need for adaptive designs that balance efficiency with personalization and user control. Future research should investigate the long-term effects of such systems on user behavior, cross-cultural differences in adoption, and the effectiveness of the QA-based approach across varying email characteristics
@@ Appendix D Technical Details of ResQ
In this section, we provide the specific implementation details and user interface used in Study 2. We developed a prototype system consisting of a Chrome extension and a backend service to enable participants to reply to emails using Gmail on a PC. The Chrome extension detected the initiation of the reply task when participants clicked the “Reply with AI” button in the Gmail reply box (see Fig. 8). Upon clicking the button, the extension extracted the email content directly from Gmail’s DOM structure using JavaScript and sent it to a backend API endpoint implemented with FastAPI  [^4] https://fastapi.tiangolo.com. The backend, hosted on an AWS EC2 instance  [^5] https://aws.amazon.com/ec2/, received the email content and forwarded it to the OpenAI API  [^6] https://platform.openai.com/docs/ to generate questions or reply suggestions. These outputs were then returned to the Chrome extension and displayed to the participant in a new reply editor. Finally, participants revise the reply suggestions and submit them back to the Gmail reply box by clicking the “Reply” button. To ensure privacy, neither the email content nor the participants’ responses were accessible to the experimenters or stored on the server.

### v2

@@ 6.2.1. Perceived Quality of the Email by Evaluators (H2)
In Fig. 6, H2 shows the results regarding the quality of the emails. The Cronbach’s Alpha of three items measuring the perceived quality of the email is $0.846$. After checking the data normality assumption with the Shapiro-Wilk test, the result of one-way repeated measures ANOVA showed that there was a significant difference in the perceived quality of the email across three conditions $(F[2,22]=9.1,p=0.001{\color[rgb]{0,0,0},\eta_{p}^{2}=0.45})$. Post-hoc analysis with Holm correction revealed that the perceived quality of the emails participants wrote in the QA-based condition was significantly higher compared to the No-AI $(t(11),p=0.005{\color[rgb]{0,0,0},d=1.21})$ condition. Thus, H2 was partially supported. The QA-based approach improved the quality of the email responses compared to the No-AI condition.
@@ 6.3.1. Perceived Impression of Participants by Evaluators (H3-a)
The results of the perceived impression of the participants rated by another group of evaluators are shown in Fig. 6 H3-a. The two items assessing participants’ impression as email senders showed high internal consistency, with a Cronbach’s Alpha of $0.946$. After checking the data normality assumption with the Shapiro-Wilk test, the result of one-way repeated measures ANOVA showed that there was a significant difference in impression of the participants as an email sender across three conditions $(F[2,22]=5.9,p=0.009{\color[rgb]{0,0,0},\eta_{p}^{2}=0.35})$. Post-hoc analysis with Holm correction revealed that participants’ impression in the QA-based condition was not significantly higher compared to both the No-AI $(t(11),p=0.058{\color[rgb]{0,0,0},d=0.79})$ and the Prompt-based $(t(11),p=0.939{\color[rgb]{0,0,0},d=0.02})$ conditions. Thus, H3-a was not supported. The QA-based approach didn’t improve the impression of participants as email senders.
@@ 6.4.1. Participants’ Email-Replying Process (RQ1)
  1. “In the QA-based condition, AI summarized key points through questions and highlighted relevant sections of the email body, which facilitated my understanding of the email and reduced my overall burden” [P10].
  2. “In the QA-based condition, I could easily obtain the desired output even without the technical skills to create prompts” [P6].
  3. “By saving the time needed to read the counterpart’s text, the psychological barrier to starting the task was lowered” [P5].
@@
  1. “Since the AI prompted me with questions at the beginning, the mental effort required to start thinking about the task was eliminated, reducing the stress associated with initiating the work” [P10].
  2. “By saving the time needed to read the counterpart’s text, the psychological barrier to starting the task was lowered” [P5].
@@ 6.4.3. Relationship between Participants and Their Counterpart (RQ3)
Participants shared differing views on how AI’s involvement affected their psychological distance from their counterparts. P2, P9, and P11 reported that the psychological distance they felt from the other person was directly related to the amount of effort they put in. Furthermore, P6 noted that “especially under QA-based condition, I barely thought about the counterpart because I only selected options to create responses.” In contrast, P8 reported that “compared to composing replies myself, using AI allowed me to create messages that left a better impression on my counterpart, which made the relationship feel closer.” These results suggested that, on the one hand, AI’s mediation can potentially increase the psychological distance between senders and receivers. On the other hand, it can also diminish the perceived distance from the sender due to effective impression management. Thus, we conducted a field study to further clarify the impact of AI on interpersonal relationships.
@@ 9.1.2. Potential Reduction in Sense of Agency and Control
While the QA-based approach enhanced users’ efficiency, our studies also revealed a potential trade-off in users’ sense of agency and control (Sec. 6.1.76.4.18.1.3). Some participants reported a decreased sense of authorship, feeling more like editors than creators of their emails. This reduction in agency may be due to the diminished amount of text input required from the user, as the AI takes a more active role in content generation. Moreover, we found that the sense of agency influenced users’ preferences for future usage. Among those participants who still maintained their sense of agency, we found that they tended to actively review and modify the AI-generated content to reflect their personal style and intentions. This suggests that even when AI intervention is substantial, users can maintain a sense of authorship by actively engaging with and refining the AI’s suggestions. To optimize users’ level of agency, adapting the degree of AI intervention in the email construction process can be helpful. For instance, by adjusting the number and type of AI-generated questions or varying the levels of AI-generated suggestions (Fu et al., 2023), ranging from word-level to message-level.
@@ 10. Conclusion
In formal email communication, users are often required to read detailed (lengthy or complex) emails. Crafting appropriate responses to such emails is time-consuming and may lead to overlooked sender requests or delayed responses, causing communication issues. Thus, we propose QA-based approach, which leverages LLM-based question generation to help users create efficient and high-quality replies by generating multiple question-answer pairs related to the received email content. To examine the comprehensive impact of the QA-based approach on both email senders and recipients, we conducted controlled and field experiments using our prototype system, ResQ. Our findings demonstrate that structuring email content into question-answer pairs improves efficiency, reduces cognitive load, and lowers barriers to initiating responses. Additionally, this approach enhances email quality and may leave a better impression on recipients. However, our findings also revealed challenges, including a potential reduction in user agency and an increased psychological distance in communication. These trade-offs emphasize the need for adaptive designs that balance efficiency with personalization and user control. Future research should investigate the long-term effects of such systems on user behavior, cross-cultural differences in adoption, and the effectiveness of the QA-based approach across varying email characteristics
@@ Appendix D Technical Details of ResQ
In this section, we provide the specific implementation details and user interface used in Study 2. We developed a prototype system consisting of a Chrome extension and a backend service to enable participants to reply to emails using Gmail on a PC. The Chrome extension detected the initiation of the reply task when participants clicked the “Reply with AI” button in the Gmail reply box (see Fig. 8). Upon clicking the button, the extension extracted the email content directly from Gmail’s DOM structure using JavaScript and sent it to a backend API endpoint implemented with FastAPI  [^4] https://fastapi.tiangolo.com. The backend, hosted on an AWS EC2 instance  [^5] https://aws.amazon.com/ec2/, received the email content and forwarded it to the OpenAI API  [^6] https://platform.openai.com/docs/ to generate questions or reply suggestions. These outputs were then returned to the Chrome extension and displayed to participant in a new reply editor. Finally, participants revise the reply suggestions and submit them back to the Gmail reply box by clicking the “Reply” button. To ensure privacy, neither the email content nor the participants’ responses were accessible to the experimenters or stored on the server.